OpenStudy (el_arrow):

find equations of the tangent plane and the normal line to the given surface at the specified point

OpenStudy (el_arrow):

xy^2z^3=8 (2,2,1)

OpenStudy (el_arrow):

i know what to do i just dont know how to get started with this one

OpenStudy (el_arrow):

@freckles could you help real quick with this problem please?

OpenStudy (daniel.ohearn1):

What's the problem, can I help?

OpenStudy (el_arrow):

its at the top

OpenStudy (daniel.ohearn1):

Do you have wolfram?

OpenStudy (daniel.ohearn1):

\[(xy ^{2z}^3 ) = 8\] Is this your problem?

OpenStudy (el_arrow):

no its \[xy ^{2}z ^{3}=8\]

OpenStudy (daniel.ohearn1):

with the point (2,2,1)

OpenStudy (el_arrow):

yes

OpenStudy (daniel.ohearn1):

What are you trying to find?

OpenStudy (el_arrow):

the equation of the tangent plane to the given surface at the specified point

OpenStudy (daniel.ohearn1):

You know the equation for the tangent plane?

OpenStudy (el_arrow):

no...thats what im looking for

OpenStudy (daniel.ohearn1):

z-z\[z - z _{0} = f _{x}(x _{0},y _{0})(x-x _{0}) + f _{y}(x _{0},y _{0})(y-y _{0})\]

OpenStudy (el_arrow):

okay but i got z^3

OpenStudy (daniel.ohearn1):

z is your point z after you solve for z and take the third root

OpenStudy (el_arrow):

what about the x and y^2

OpenStudy (irishboy123):

the cheater's way !! for \(xy^2z^3=8\) make a level surface \(f = xy^2z^3\) and then compute \(\nabla f = <y^2z^3, 2xyz^3, 3xy^2z^2> \) at \( (2,2,1)\) for the normal vector

OpenStudy (el_arrow):

okay and what is the next step? @IrishBoy123

OpenStudy (irishboy123):

well that's it all pretty much done !! put (x,y,z) = (2,2,1) into that \(\nabla f\) vector and that is the normal \(\vec n\) the equation for the plane is then \(<x,y,z>.\vec n = const\) and you have a point on the plane, ie (2,2,1), to get that constant hope that helps!

OpenStudy (el_arrow):

thanks @IrishBoy123 couldnt have done it without your help!